OpenStudy (anonymous):

find all values of k such that x^2-kx+9 = 0 has two imaginary roots

OpenStudy (mathmath333):

do u know discriminant of this equaation

OpenStudy (anonymous):

6

OpenStudy (anonymous):

but thats not right apparently

OpenStudy (mathmath333):

formula for discriminant \(\large \color{black}{\begin{align} if\quad ax^2+bx+c=0\hspace{.33em}\\~\\ then\quad D=b^2-4ac\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

can u find \(D\) now?

OpenStudy (anonymous):

-k^2-36

OpenStudy (mathmath333):

some fault there , note that \(\large b=-k\) here

OpenStudy (anonymous):

yeah and i said (-k)^2

OpenStudy (mathmath333):

\(-k^2\neq (-k)^2\)

OpenStudy (mathmath333):

\((-k)^2=k^2\)

OpenStudy (anonymous):

ok

OpenStudy (mathmath333):

ok the discriminant property says that the equation has imaginary roots if \(\large \color{black}{\begin{align} D<0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

yes

OpenStudy (mathmath333):

so \(\large \color{black}{\begin{align} k^2-36<0\hspace{.33em}\\~\\ \end{align}}\) can u solve this

OpenStudy (anonymous):

k< 6

OpenStudy (mathmath333):

that's not complete, use the property that \(\large \color{black}{\begin{align} a^2-b^2=(a-b)(a+b)\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

can you tell me what you got

OpenStudy (mathmath333):

i got \(\large \color{black}{\begin{align} k^2-36<0\hspace{.33em}\\~\\ k^2-6^2<0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

isnt it k> -6

OpenStudy (mathmath333):

now we have \(\large \color{black}{\begin{align} k^2-6^2<0\hspace{.33em}\\~\\ (k-6)(k+6)<0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

so we have \(\large \color{black}{\begin{align} k<6\hspace{.33em}\\~\\ and \hspace{.33em}\\~\\ k>-6\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

so \(\large \color{black}{\begin{align} -6<k<6\hspace{.33em}\\~\\ \end{align}}\)